# A Deconfounding Approach to Climate Model Bias Correction

### v1

@@ Introduction
$\mathbf{E}[Y(\bar{\mathbf{a}}_{\geq t})\mid\bar{\mathbf{A}}_{t-1},\mathbf{X}_{t},\bar{\mathbf{Z}}_{t}]=\mathbf{E}[Y\mid\bar{\mathbf{a}}_{\geq t},\bar{\mathbf{A}}_{t-1},\mathbf{X}_{t},\bar{\mathbf{Z}}_{t}].$
This equation implies that the potential outcome can be represented by $\mathbf{E}[Y\mid\bar{a}_{\geq t},\bar{\mathbf{A}}_{t-1},\bar{\mathbf{X}}_{t},\bar{\mathbf{Z}}_{t}]$. Therefore, our estimate of the potential outcome is unbiased given the previous treatments, current covariates, and latent variables. This unbiasedness is crucial for making valid inferences about causal effects.
Then, we will show how to implement in practical. The joint distribution of all collected T time steps in our enhanced Deconfounding BC Factor Model can be formulated as follows:
$p(\theta_{1:k}^{G},\bar{\mathbf{x}}_{T}^{G},\bar{\mathbf{z}}_{T},\bar{\mathbf{a}}_{T}^{G})=$
$p(\theta_{1:k}^{G})p(\bar{\mathbf{x}}_{T}^{G})\prod_{t=1}^{T}p(\mathbf{z}_{t}\mid\bar{\mathbf{h}}_{t-1})\prod_{j=1}^{k}p(\mathbf{a}_{tj}^{G}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{G},\theta_{j}^{G}),$
$p(\theta_{1:k}^{O},\bar{\mathbf{x}}_{T}^{O},\bar{\mathbf{z}}_{T},\bar{\mathbf{a}}_{T}^{O})=$
$p(\theta_{1:k}^{O})p(\bar{\mathbf{x}}_{T}^{O})\prod_{t=1}^{T}p(\mathbf{z}_{t}\mid\bar{\mathbf{h}}_{t-1})\prod_{j=1}^{k}p(\mathbf{a}_{tj}^{O}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{O},\theta_{j}^{O}),$
@@
In practical applications, as shown figure 2, we need to build a model to get $\mathbf{Z}$ with predicting treatments as the constraint. To manage time-varying treatments $\mathbf{A}^{G},\mathbf{A}^{O}$ , we follow a pragmatic approach utilizing a recurrent neural network (RNN) with multitask output for implementation.
@@
$\mathbf{Z}_{1}=\text{RNN}(\mathbf{L}),$
$\quad\mathbf{Z}_{t}=\text{RNN}(\mathbf{Z}_{t-1},\mathbf{X}_{t-1}^{G},\mathbf{A}_{t-1}^{G},\mathbf{X}_{t-1}^{O},\mathbf{A}_{t-1}^{O},\mathbf{L}),$
@@
$A_{tj}^{G}=\text{FC}(\mathbf{X}_{t}^{G},\mathbf{Z_{t}},\theta_{j}^{G}),$
$A_{tj}^{O}=\text{FC}(\mathbf{X}_{t}^{O},\mathbf{Z}_{t},\theta_{j}^{O}),$
for all $j$ and $t$, where $\theta_{j}^{G}$ and $\theta_{j}^{O}$ are the parameters of the fully connected layers. Using this as multitask output, treatments is conditionally independent given $\mathbf{X_{t},Z_{t}}$ for both sources, this can be ensured by inferring each treatment $A^{G}$ or $A^{O}$ as different tasks with only $\mathbf{X}_{t}^{G},\mathbf{Z_{t}}$ or $\mathbf{X}_{t}^{O},\mathbf{Z_{t}}$, if we condition on $\mathbf{X}_{t}^{G},\mathbf{Z_{t}}$ or $\mathbf{X}_{t}^{O},\mathbf{Z_{t}}$, they are conditionally independent.
Figure 4: The overall process of Deconfounding Bias Correction. The process begins with the Deconfounding BC factor model, which extracts a latent variable $\mathbf{Z}$ from both Global Climate Model (GCM) data and observational (OBS) data. This latent variable represents unobserved confounders that affect both datasets. The iTransformer then utilizes this latent variable $\mathbf{Z}$ to generate correction terms for the GCM predictions, adjusting for bias and producing more accurate future climate predictions. This general process ensures that the GCM outputs are corrected for taking into account unobserved confounders.
Training
@@
After obtaining hidden multi-cause confounders $\mathbf{Z}$ from both GCM and observational data in one area. In our problem formulation, we aim to correct the bias between $Y^{G}$ and $Y^{O}$ while considering unobserved confounders. A key question is how to utilize this information for bias correction. Advanced time series forecasting models such as iTransformer (Liu et al. 2024) have demonstrated significant potential in capturing temporal information, and integrating these models into our bias correction approach promises substantial improvements.
@@
We propose to use probability model for bias correction. As Figure 4 shown, we build a prediction model based on the observational data and the multi-cause confounder Z to predict future precipitation.
The probability model employed in is to obtain predicted $\Delta{Y}$. The probability model can theoretically be any time series prediction model capable of generating samples from $P_{\theta}(\Delta{Y}_{t+k}|\mathbf{A}^{G},\mathbf{A}^{O},\mathbf{Z})$, possessing an explicit likelihood form, and utilizing the constructed training data for this purpose. For instance, a linear regression model may serve as the forecasting model. However, due to its inherent assumption of independence among the elements of $\Delta{Y}_{t+k}$, it is likely to result in a poor fit.
To perform well in such a forecasting task. We choose the state of art forecasting model iTransformer (Liu et al. 2024). The reason why we choose this model is: first, its performance in long-term forecasting is great. second, it is a simple model that is an inverted transformer designed for time series. More details about iTransformer can be found in the Appendix.
After we get the $\Delta{Y}_{t+k}$, We can obtain our corrected GCM output with:
$\{\hat{Y}_{t+1}^{G},\hat{Y}_{t+2}^{G},\dots,\hat{Y}_{t+k}^{G}\}=\{Y_{t+1}^{G}+\Delta Y_{t+1},$
$Y_{t+2}^{G}+\Delta Y_{t+2},\dots,Y_{t+k}^{G}+\Delta Y_{t+k}\}$
Experiments and results
@@
The objectives of the experiments are as follows: 1) Use synthetic data sets to evaluate the correctness of the latent confounder captured by our deconfounding method. 2) Build a correction model to correct precipitation predictions made by the climate model. Implementation code will be given later.
Experiments on Synthetic Data
To assess the effectiveness of our deconfounding method proposed in this paper, we conducted experiments using synthetic data in which we use predictive analysis to assess the influence of hidden confounders. Validation with real-world data is challenging because the true impact of hidden confounding cannot be precisely determined (Wang and Blei 2019).
Simulated dataset
To maintain the generality of the simulation process, we propose generating a dataset using a two-source autoregressive model based on the causal summary graph shown in Figure 1(c). At each timestep $t$, we simulate time-varying covariates $X_{t}$ from two sources ($X_{t}^{\text{source1}}$ and $X_{t}^{\text{source2}}$), along with a multi-cause hidden confounder $Z_{t}$. A detailed description of the data generation process is provided in the Appendix.
We create datasets that include 500 locations, 3650 timesteps, and $k=3$ covariates and treatments. To introduce time dependencies, we assign $p=5$. Each dataset is divided into an 80/10/10 split for training, validation, and testing purposes, respectively.
Our method’s effectiveness relies on two key points. First, the inferred latent variable $\mathbf{Z}$ must, along with $\mathbf{X}$, accurately predict the treatment $\mathbf{A}$. This ensures that $\mathbf{A}$ is conditionally independent given $\mathbf{Z}$ and $\mathbf{X}$. Second, and most importantly, our method aims to learn the latent confounder $\mathbf{Z}$. If the inferred $\mathbf{Z}$ closely matches or has a similar distribution to the simulation-generated $\mathbf{Z}$, it demonstrates that our method successfully captures the true latent confounders.
The theory posits that using inferred latent variables as substitutes for hidden confounders can yield unbiased outcomes, leveraging the factor model’s ability to accurately capture the distribution of the assigned causes. If the multi-layer output aligns well with the generated treatment, we can assert that our method effectively captures the treatment and ensures these treatments are conditionally independent when conditioned on covariates and confounders, as previously discussed.
To validate this, we performed a predictive check. The mean squared error (MSE) for the assigned treatment is 0.06157, indicating the effectiveness of our method in capturing treatments. Additionally, we compared the inferred $\mathbf{Z}$ with the simulation-generated $\mathbf{Z}$. The resulting MSE is 0.00187. This further demonstrates the effectiveness of our method in learning latent confounders.
Case study: South Australia
@@
Figure 5: Study Area: South Australia (Latitude: -35.0° to -28.0°, Longitude: 129.0° to 141.0°)
@@
For observational reference data, NCEP-NCAR Reanalysis 1222NCEP-NCAR data portal: https://psl.noaa.gov/data/gridded/data. ncep.reanalysis.html, provided by the National Oceanic and Atmospheric Administration (NOAA), is utilized. The dataset encompasses the same variables as the IPSL model (e.g., tmax & prate). NCEP-NCAR reanalysis data is available at a monthly frequency, covering the period from 1948 to 2014, with a global resolution of 2.5 degrees in both the latitudinal and longitudinal directions, and has not been re-gridded. The nearest geographical point to South Australia was selected.
@@
This research concentrates on bias correction of the precipitation rate (prate) to improve the accuracy of precipitation predictions.
Figure 6: QQ plots for South Australia, 1993-2014, comparing observed precipitation (X-axis) with precipitation estimated by bias correction (BC) methods (Y-axis). The closer the alignment to the red line, the better the performance.
Data Preprocessing
@@
We extracted and converted climate data from 1948 to 2014 for the South Australia region as shown in Figure 5. The data, originally in NetCDF format, were transformed into CSV files for further processing. Subsequently, we split the data into a training period (1948-1992) and a testing period (1993-2014). The IPSL dataset contains 30 initial conditions; to generalize our model, we selected 15 initial conditions for the experiment. In this paper, we present results from three of these initial conditions(r5i1p1f1, r6i1p1f1 and r7i1p1f1) along with their average results. Complete results can be found in the Appendix.
Experiment settings
@@
To evaluate our proposed method, we compare our methods with several bias correction baseline methods, including linear scaling (Teutschbein and Seibert 2012b), variance scaling (Teutschbein and Seibert 2012b), quantile mapping (Cannon, Sobie, and Murdock 2015), quantile delta mapping (Tong et al. 2021) and Temporal BC (Nivron et al. 2024) methods. Implementation details of baselines are introduced in Appendix. The metric we are using to compare the performance is the Mean Squared Error (MSE) and the Mean Absolute Error (MAE).
Results
@@
Table 2 illustrates the three-step-ahead predictions using our deconfounding bias correction method on the dataset with different GCM initial settings. After applying the deconfounding step and augmenting the dataset with substitutes for the hidden confounders, we evaluated the performance of various bias correction methods using MSE and MAE. The results in Table 2 support the effectiveness of the Deconfounding BC method, which consistently achieves the lowest MSE and MAE across all experimental conditions. This indicates its superior performance in reducing prediction errors and aligning closely with the observed data.
@@
The QQ plots (Figure 6) reveal that among the evaluated bias correction methods, Deconfounding BC shows the closest alignment with the red dashed line (y=x) across the entire range of observed precipitation values. This indicates that Deconfounding BC is the most effective in accurately capturing the observed precipitation levels, outperforming other methods in reducing biases and providing reliable predictions.
Figure 7: Box plots of monthly precipitation values for South Australia using different bias correction methods.
Figure 7 presents box plots of monthly precipitation values for South Australia using different bias correction methods, with ‘obsp‘ representing the observed (real) values. Among the methods compared, Deconfounding BC stands out for its effectiveness in bis correction.. The box plot for Deconfounding BC shows a median and distribution range that closely match the observed values, indicating that this method captures the variability and distribution of real precipitation data more accurately than other methods.
Ablation Study
@@
To demonstrate that the latent confounder learned by our Deconfounding BC factor model contains essential information, we compared our correction model’s results with and without the latent confounder, as shown in Table 3. The performance improved significantly with the inclusion of the hidden confounder.
Conclusion
@@
In this paper, we proposed the deconfounding bias correction method, with the presence of multi-cause confounders. By integrating climate bias correction techniques with advancements in causality based time series deconfoudning method, our approach offers a novel perspective for future studies. This highlights the importance of not always assuming that all variables are observed.

### v2

@@ 3.1.1 Deconfounding BC Factor Model
${E}[Y(\bar{a}_{\geq t})\mid\bar{\mathbf{A}}_{t-1},\bar{\mathbf{X}}_{t},\bar{\mathbf{Z}}_{t}]={E}[Y\mid\bar{a}_{\geq t},\bar{\mathbf{A}}_{t-1},\bar{\mathbf{X}}_{t},\bar{\mathbf{Z}}_{t}].$ (10)
This equation implies that the potential outcome can be represented by ${E}[Y\mid\bar{a}_{\geq t},\bar{\mathbf{A}}_{t-1},\bar{\mathbf{X}}_{t},\bar{\mathbf{Z}}_{t}]$. Therefore, our estimate of the potential outcome is unbiased given the previous treatments, current covariates, and latent variables. This unbiasedness is crucial for making valid inferences about causal effects.
Then, we will show how to implement in practical based on factorization Formula (3) and (4). The joint distribution of all collected T time steps in our enhanced Deconfounding BC Factor Model can be formulated as follows:
$p(\theta_{1:k}^{G},\bar{\mathbf{x}}_{T}^{G},\bar{\mathbf{z}}_{T},\bar{\mathbf{a}}_{T}^{G})=$ $\,p(\theta_{1:k}^{G})p(\bar{\mathbf{x}}_{T}^{G})\prod_{t=1}^{T}p(\mathbf{z}_{t}\mid\bar{\mathbf{h}}_{t-1})$ (11)
×$\times\prod_{j=1}^{k}p(\mathbf{a}_{tj}^{G}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{G},\theta_{j}^{G}),$
The same formulation holds for the GCM source by replacing $O$ with $G$.
Figure 4: Workflow of the Deconfounding BC factor model(model architecture refers to Figure 3) using simulation data: the GCM Outputs and observational data are combined to estimate the unobserved confounder $Z$ through the Deconfounding BC factor model. The iTransformer utilizes $Z$ to compute corrections ($\Delta Y_{t+1},\ldots,\Delta Y_{t+k}$) and adds them to the GCM predictions ($Y_{t+1}^{G},\ldots,Y_{t+k}^{G}$) to generate the final corrected precipitation predictions.
@@
In practical applications, as shown Figure 3, we need to build a model to obtain $\mathbf{Z}$ with predicting treatments as the constraint. To manage time-varying treatments $\mathbf{A}^{G},\mathbf{A}^{O}$ , we follow a pragmatic approach by utilizing a recurrent neural network (RNN) with multitask output for implementation.
@@
$\mathbf{Z}_{1}=\text{RNN}(\mathbf{L}),$ (12)
$\quad\mathbf{Z}_{t}=\text{RNN}(\mathbf{Z}_{t-1},\mathbf{X}_{t-1}^{G},\mathbf{A}_{t-1}^{G},\mathbf{X}_{t-1}^{O},\mathbf{A}_{t-1}^{O},\mathbf{L}),$ (13)
@@
$A_{tj}^{G}=\text{FC}(\mathbf{X}_{t}^{G},\mathbf{Z_{t}},\theta_{j}^{G}),$ (14)
$A_{tj}^{O}=\text{FC}(\mathbf{X}_{t}^{O},\mathbf{Z}_{t},\theta_{j}^{O}),$ (15)
for all $j$ and $t$, where $\theta_{j}^{G}$ and $\theta_{j}^{O}$ are the parameters of the fully connected layers. Using these as multitask output, treatments is conditionally independent given $\mathbf{X_{t},Z_{t}}$ for both sources, this can be ensured by inferring each treatment $A^{G}$ or $A^{O}$ as different tasks with only $\mathbf{X}_{t}^{G},\mathbf{Z_{t}}$ or $\mathbf{X}_{t}^{O},\mathbf{Z_{t}}$, if we condition on $\mathbf{X}_{t}^{G},\mathbf{Z_{t}}$ or $\mathbf{X}_{t}^{O},\mathbf{Z_{t}}$, they are conditionally independent.
3.1.2 Training
@@ 3.2 Correction
After obtaining hidden multi-cause confounders $\mathbf{Z}$ from both GCM and observational data in one area. In our problem formulation, we aim to correct the bias between $Y^{G}$ and $Y^{O}$ while considering unobserved confounders. A key question is how to utilize this information for bias correction. Advanced time series forecasting models such as iTransformer Liu et al. (2024) have demonstrated significant potential in capturing temporal information, and integrating these models into our bias correction approach promises substantial improvements. Compared to models like TimeMixer and PatchTST, iTransformer balances complexity and performance.
@@
We propose to use a probability model for bias correction. As shown in Figure 4, we build a prediction model based on the observational data and the multi-cause confounder Z to predict future precipitation.
The probability model employed in our framework is designed to generate predictions of $\Delta Y$, the correction term for the GCM output. Formally, the model learns to approximate the conditional distribution: $P_{\theta}(\Delta Y_{t+k}\mid\mathbf{A}^{G},\mathbf{A}^{O},\mathbf{Z}),$ where $\mathbf{A}^{G}$ and $\mathbf{A}^{O}$ are the current covariates from GCM and observations, and $\mathbf{Z}$ is the inferred latent confounder.
To achieve high performance in forecasting tasks involving complex temporal dependencies, we choose to use iTransformer Liu et al. (2024) as the prediction model. iTransformer is a lightweight and effective transformer variant specifically designed for long-sequence time series forecasting, offering a strong balance between expressiveness and efficiency.
After we obtain the $\Delta{Y}_{t+k}$, we can derive our corrected GCM output with:
$\begin{split}\{\hat{Y}_{t+1}^{G},\hat{Y}_{t+2}^{G},\dots,\hat{Y}_{t+k}^{G}\}=\\ \{Y_{t+1}^{G}+\Delta Y_{t+1},Y_{t+2}^{G}+\Delta Y_{t+2},\dots,Y_{t+k}^{G}+\Delta Y_{t+k}\}\end{split}$ (16)
4 Experiments And Results
@@
The objectives of the experiments are as follows: 1) Use synthetic data sets to evaluate the correctness of the latent confounder captured by our deconfounding method. 2) Build a correction model to correct precipitation predictions made by the climate model. Implementation code is attached.
4.1 Experiments On Synthetic Data
To assess the effectiveness of our deconfounding method, we conduct controlled experiments on synthetic data, since the true influence of hidden confounders cannot be measured directly in real‐world settings Wang and Blei (2019). We generate data from a two‐source autoregressive model based on the causal summary graph in Figure 1(c). At each timestep $t$, we sample time‐varying covariates $X_{t}^{\text{source1}}$ and $X_{t}^{\text{source2}}$, together with a multi‐cause latent confounder $Z_{t}$. We simulate $N=500$ locations over $T=3650$ timesteps, with $k=3$ covariates and treatments and autoregressive order $p=5$. The resulting dataset is split into 80% training, 10% validation, and 10% testing.
4.1.1 Simulated Dataset
(a) Time series plot
(b) QQ plot
(c) Box plot
Figure 5: Comparison of bias correction methods for precipitation using simulation data: (Left) Time series alignment with observed data, demonstrating the performance of Deconfounding BC and other methods; (Center) QQ plot comparing predicted rainfall against observed values, highlighting Deconfounding BC’s closer alignment with the ideal line; (Right) Box plot of rainfall distributions across bias correction methods, illustrating variability and alignment with observed data.
Our evaluation focuses on two properties. First, the inferred latent variable $\mathbf{Z}$, together with the observed covariates $\mathbf{X}$, should accurately predict the treatment vector $\mathbf{A}=[A_{1},\ldots,A_{k}]$. Concretely, we model
$p(\mathbf{A}\mid\mathbf{Z},\mathbf{X})=\prod_{j=1}^{k}p\bigl{(}A_{j}\mid\mathbf{Z},\mathbf{X}\bigr{)}$
Second, the learned $\mathbf{Z}$ should match the ground‐truth confounders used in simulation. A low mean squared error (MSE) between inferred and true $\mathbf{Z}$ indicates successful recovery of the hidden structure. The details for data generation can be found in our GitHub repository.
Figure 5 summarizes our predictive checks. The Deconfounding BC method yields an MSE of $0.06157$ when predicting the treatment vector, demonstrating strong alignment between model predictions and the simulated treatments. Comparing the inferred $\mathbf{Z}$ against the true latent confounders gives an MSE of $0.00187$, confirming that our factor‐model approach accurately captures the hidden variables. Together, these results validate that the treatments become conditionally independent given the inferred confounders and covariates, and that the learned latent representation faithfully recovers the ground‐truth confounding signal.
4.2 Case Study: South Australia
@@
For observational reference data, NCEP-NCAR Reanalysis 1333NCEP-NCAR data portal: https://psl.noaa.gov/data/gridded/data. ncep.reanalysis.html, provided by the National Oceanic and Atmospheric Administration (NOAA), is utilized. The dataset encompasses the same variables as the IPSL model (e.g., tmax & prate). NCEP-NCAR reanalysis data is available at a monthly frequency, covering the period from 1948 to 2014, with a global resolution of 2.5 degrees in both the latitudinal and longitudinal directions, and has not been re-gridded. The nearest geographical point to South Australia was selected.
@@
This research focuses on bias correction of the precipitation rate to improve the accuracy of precipitation predictions.
4.2.1 Data Preprocessing
@@
We extracted and converted climate data from 1948 to 2014 for the South Australia region. The data, originally in NetCDF format, were transformed into CSV files for further processing. Subsequently, we split the data into a training period (1948-1992) and a testing period (1993-2014). The IPSL dataset contains 30 initial conditions; to generalize our model, we selected 15 initial conditions for the experiment. In this paper, we present results from three of these initial conditions(r5i1p1f1, r6i1p1f1 and r7i1p1f1) along with their average results.
4.2.2 Experiment Settings
@@
To evaluate our proposed method, we compare our method with several bias correction baseline methods, including linear scaling Teutschbein and Seibert (2012), variance scaling Teutschbein and Seibert (2012), quantile mapping Cannon et al. (2015), quantile delta mapping Tong et al. (2021) and Temporal BC Nivron et al. (2024) methods. The metric we are using to compare the performance is the Mean Squared Error (MSE) and the Mean Absolute Error (MAE).
4.3 Results
@@
Table 2 illustrates the three-step-ahead predictions using our deconfounding bias correction method on the dataset with different GCM initial settings. After applying the deconfounding step and augmenting the dataset with substitutes for the hidden confounders, we evaluated the performance of various bias correction methods using MSE and MAE. The results in Table 2 support the effectiveness of the Deconfounding BC method, which consistently achieves the lowest MSE and MAE across all experimental conditions. This indicates its superior performance in reducing prediction errors and aligning closely with the observed data. Figure 6 further visualizes the daily precipitation trends from IPSL, observations, and various correction methods, clearly highlighting the superior alignment achieved by Deconfounding BC.
Figure 6: Comparison of daily precipitation rates from IPSL , observation and various bias correction methods over a year, highlighting the performance of Deconfounding BC in aligning with observed trends.
@@
The QQ plots (Figure 7) reveal that among the evaluated bias correction methods, Deconfounding BC shows the closest alignment with the red dashed line (y=x) across the entire range of observed precipitation values. This indicates that Deconfounding BC is the best in accurately capturing the observed precipitation levels, outperforming other methods in reducing biases and providing reliable predictions.
Figure 7: This figure shows Quantile-Quantile (QQ) plots comparing observed and model precipitation quantiles using various bias correction methods. The left panel highlights the performance of different bias correction methods, with a focus on how closely they align with the 1:1 line, which indicates perfect agreement. The right panel illustrates overall agreement across all quantiles, providing a broader comparison of the methods.
@@
To demonstrate that the latent confounder learned by our Deconfounding BC factor model contains essential information, we did ablation study to compare our correction model’s results with and without the latent confounder, as shown in Table 3. The performance improved with the inclusion of the hidden confounder.
5 Conclusion
@@
In this paper, we propose the deconfounding bias correction method for multi-cause confounders. By integrating climate bias correction techniques with causality-based time series deconfounding, our approach provides a novel perspective for future studies, emphasizing the importance of not assuming all variables are observed.
Acknowledgments
This work was supported by the ARC Discovery Project DP230101122 and the University of South Australia Research Training Program (RTP) Scholarship. We gratefully acknowledge the continued support from the CSIRO Environment Research Unit and the Data61 Business Unit.